# Automatically Labeling \$200B Life-Saving Datasets: A Large Clinical Trial Outcome Benchmark

### v2

@@ Discussion
The CTO dataset provides a comprehensive view of clinical trial data, capturing key information across multiple dimensions for around  125K drug and biologics trials. Figure 4A, B, and D provides an overview of the dataset’s key statistics, highlighting its scale and diversity. Figure 4A illustrates the annual distribution of completed clinical trials, and it shows a steady increase over the years, reflecting the growing scale of global clinical research. In Figure 4B, we show the frequency distribution of the top 30 medical conditions within the dataset, where Neoplasms are the most represented condition. The phase distribution of trials in the CTO dataset is shown in Figure 4D. Note that we ensured that our set of manual labels accurately reflects the overall clinical trial distribution as shown in Figure 2A, B and C.
@@
To assess the agreement between the labeling functions, we calculated the percentage of agreement for trials that fall within the intersections of their coverage, as shown in Figure 4C. The heatmap highlights the agreement scores only for the important labeling functions in the dataset. Each cell in the heatmap represents the agreement score between a pair of labeling functions. It is important to note that each labeling function does not cover all trials with weak labels. Therefore, while calculating the agreement score, we only considered the trials common to both labeling functions. For example, there are no common trials with weak labels between the news headlines and trial status labeling functions (represented as a black cell). This is because the trial status labeling function assigns a value of -1 (indicating an inability to decide the trial outcome) to trials with a ‘completed’ status, while news extraction is limited to ‘completed’ trials to optimize resource usage associated with SerpAPI.
@@
Another noteworthy observation is the low agreement between p-values and trial status. This discrepancy arises because most trial status-based failure labels correspond to trials marked as ‘terminated,’ ‘withdrawn,’ or ‘suspended,’ often due to sponsor or business decisions rather than scientific outcomes. Interestingly, many of these trials report statistically significant p-values ($<0.05$) for some of their primary endpoints but are stopped for reasons unrelated to efficacy. For instance, our analysis identified approximately 1,200 trials marked as ‘terminated’ that reported primary outcome p-values below 0.05. Specific examples include NCT04364763, which was terminated due to enrollment feasibility issues, and NCT01374451, which was stopped for failing to meet the primary endpoint of progression-free survival. These cases highlight the complex interplay between trial outcomes and external factors.
@@ Conclusion
We successfully created the first large-scale, publicly available knowledge base of clinical trial outcomes that automatically produces useful outcome labels. We additionally release a set of manually annotated outcome labels using human experts. Since we are the first to publicly release our trial annotation pipeline along with a corresponding knowledge base, we hope to democratize trial outcome prediction, enabling faster, more transparent trial design. Our dataset of recent trials, weak labels, and knowledge base is available at https://chufangao.github.io/CTOD/.

### v3

@@ Discussion
The CTO dataset provides a comprehensive view of clinical trial data, capturing key information across multiple dimensions for around  125K drug and biologics trials. Figure 4A, B, and D provide an overview of the dataset’s key statistics, highlighting its scale and diversity. Figure 4A illustrates the annual distribution of completed clinical trials, and it shows a steady increase over the years, reflecting the growing scale of global clinical research. In Figure 4B, we show the frequency distribution of the top 30 medical conditions within the dataset, where Neoplasms are the most represented condition. The phase distribution of trials in the CTO dataset is shown in Figure 4D. Note that we ensured that our set of manual labels accurately reflects the overall clinical trial distribution as shown in Figure 2A, B, and C.
@@
To assess the agreement between the labeling functions, we calculated the percentage of agreement for trials that fall within the intersections of their coverage, as shown in Figure 4C. The heatmap highlights the agreement scores only for the important labeling functions in the dataset. Each cell in the heatmap represents the agreement score between a pair of labeling functions. It is important to note that each labeling function does not cover all trials with weak labels. Therefore, we only considered the trials common to both labeling functions while calculating the agreement score. Because of this, there are no common trials with weak labels between the news headlines and trial status labeling functions (represented as a black cell). This is because the trial status labeling function assigns a value of -1 (indicating an inability to decide the trial outcome) to trials with a ‘completed’ status, while news extraction is limited to ‘completed’ trials to optimize resource usage associated with SerpAPI.
@@
Another noteworthy observation is the low agreement between p-values and trial status. This discrepancy arises because most trial status-based failure labels correspond to trials marked as ‘terminated,’ ‘withdrawn,’ or ‘suspended,’ often due to sponsor or business decisions rather than any trial outcome. Interestingly, some of these trials report statistically significant outcomes (defined as p-values($<0.05$) for at least one of their primary endpoints but were stopped for reasons unrelated to efficacy. For instance, our analysis identified approximately 1,200 trials marked as “terminated” that reported at least one statistically significant primary outcome. Specific examples include NCT04364763, which was terminated due to enrollment feasibility issues, and NCT01374451, which was stopped for failing to meet the primary endpoint of progression-free survival. These cases highlight the complex interplay between trial outcomes and external factors.
@@ Conclusion
We present the first large-scale, publicly available knowledge base for clinical trial outcomes, providing automatically generated outcome labels alongside a manually curated dataset annotated by human experts. By publicly releasing our trial annotation pipeline and the corresponding knowledge base, we aim to democratize trial outcome prediction, fostering more transparent and data-driven clinical trial design. Our dataset, including recent trials, weak labels, and the full knowledge base, is available at https://chufangao.github.io/CTOD/, with regular updates to support ongoing research and innovation.